# Search for cosmic rays in GRANDProto300

### v1

@@ Abstract
GRANDProto300 (GP300) is a prototype array of the GRAND experiment, designed to validate the technique of autonomous radio-detection of astroparticles by detecting cosmic rays with energies between 1017-1018.5 eV. This observation will further enable the study of the Galactic-to-extragalactic source transition region. Since November 2024, 46 out of 300 antennas have been operational and collecting data stably. We present our cosmic-ray search pipeline, which involves several filtering steps: (1) coincidence search for signals triggering multiple antennas within a time window, (2) directional reconstruction of events, (3) exclusion of clustered (in time and space) noise events, (4) polarization cut, (5) selection based on the size of the footprint, and (6) other more arbitrary cuts in this preliminary stage, including visual cuts. The efficiency of the pipeline is evaluated and applied to the first batch of data, yielding a set of cosmic-ray candidate events, which we present.
@@ 2 Experimental setup and dataset
GP300 consists of 300 antennas spread over an area of 200 km2 situated in the Gobi desert, near Dunhuang, China (lying at N, E). As of July 2025, 65 antennas with an infill step of 577 m have been deployed and are operational, collecting data stably [2]. Each Detection Unit (DU) operates with a sampling frequency of 500 MHz and an analog-to-digital (ADC) converter digitizes signals across three distinct channels, each constructed from a dipole arm of the antenna, along a direction ( channel along North-South, channel along West-East, and channel along the vertical axis). The ADC data, recorded over a duration of 2 s in a frequency range of 50 to 200 MHz, is then converted to voltage traces via the GRANDlib software [5]. The GRAND experiment handles data in three formats: Monitoring Data (MD), continuous monitoring of the background radio environments to identify and filter out noise; Unit Data (UD), which records signals from individual DUs that pass a local trigger based on 6 parameters, including two amplitude thresholds set at 5 and 3 sigma above background noise, ensuring that only significant events are captured [6, 7]; Coincidence Data (CD), data from multiple (4) antennas triggered within a 10 s time window, providing a more accurate identification and reconstruction of the signal. The collected data is processed via the GRANDlib software [5].
@@ 2.1 Data processing: cosmic-ray search pipeline
The pipeline for cosmic-ray search is directly applied to CD events. They are subject to two types of noise: transient noise and stationary background noise. Transient noise induces events and is targeted for batch removal due to its high occurrence; for example, 70 % of all events originate from an electric transformer in the northeast. Stationary background noise, on the other hand, appears within the baseline of the trace but does not independently trigger DUs. For this pipeline, we aim to adopt a conservative approach, prioritizing the reduction of false positives over false negatives in our cosmic-ray identification process.
@@ 2.2 Clustering Cut
The GP300 dataset is affected by interference from commercial aircraft hovering near the array and regular emissions from an electric transformer located northeast of the site. This interference results in batches of noise events that can be excluded from the analysis. Given that cosmic rays produce short radio pulses ( ns) and thus cannot result in multiple CD, a clustering algorithm is well-suited for this task.
The algorithm is based on the spatial proximity of events using Planar Wave Front (PWF) arrival direction reconstruction, following the method developed in [8]. Multiple events occurring within the same angular and time window are designated as a cluster and consequently excluded from the dataset. The angular distance between events and is defined as the minimum between and . The temporal distance is defined as . Fig. 2 (left) presents the distribution of angular and time distances for consecutive events in our dataset. By varying the angular and time cut parameters below which the events are suppressed, the figure demonstrates the extent of data reduction achieved with different parameter choices. In this study, a conservative approach is adopted, and the parameters of and s are retained, resulting in a cut of 79 % of the CD data, discriminated against as background noise, as seen in Fig. 5 (right).
@@ 2.4 Systematic quality cuts
To ensure high-quality events, several additional quality cuts are implemented in the cosmic-ray search pipeline. They are independently applied to the data, and the order of the processes does not matter. We detail below all the additional cuts performed on the data. The performances of each cut, in terms of percentage of background events removed are presented in the right-hand table in Fig. 5. The cuts based on the antenna number (footprint size) and zenith angle (inclination of the shower) are built on physical grounds: GP300 focuses on specific cosmic-ray parameter ranges, that favor specific ranges in footprint size and zenith angle. The PWF error cut is more artificial as poorly reconstructed events could still be cosmic rays. However, in the current stage, events that cannot be reconstructed consistently with this PWF procedure are not considered (an error based on spherical wavefront is being implemented and tested). The last two cuts (SNR and RMS) are implemented at this stage of data commissioning in order to simplify noise discrimination and help clean events emerge from the sample. They are however to be removed and/or refined for the next stage of the pipeline development, to process the next batch of data with the 65 operational antennas.
@@
Zenith. Given its antenna sparsity, GP300 specializes in detecting highly inclined air showers with zenith angle . The final cut removes events with a reconstructed arrival direction in zenith angle. In addition, reconstruction of near-horizon event is difficult, hence we remove as well events with zenith angle . Results are shown in Fig. 5 (right).
@@ 2.5 Manual quality cuts
Trace visual cut. This visual cut is based on the shape of the signal trace in each antenna. Cosmic rays are characterized by short-duration pulses and low-frequency signals.
@@
After passing all cuts, the candidates are reconstructed using three different methods: lateral distribution function [12] from deconvolved electric field reconstruction [3], angular distribution function (ADF) directly from voltage traces [13, 14], and Graph Neural Networks directly from voltage traces [15]. The majority of the reconstructions confirm the detection of cosmic rays in the expected energy range. These reconstruction methods serve as a basis for an additional quality cut, using multiple criteria: DUs with reconstructed electric fields ( [12] and [13]), a reconstructed energy eV, and a chi-squared value associated with the ADF method lower than 25, as described in [13].
@@ 3 First batch of cosmic-ray candidates
Cut Parameters | Cut efficiency | |
Clustering | s | 79 % |
Polarization | 22 % | |
Nb of DUs | 50 % | |
Zenith | 56 % | |
PWF error | 31 % | |
SNR | 58 % | |
RMS | 85 % | |

### v2

@@ Abstract
GRANDProto300 (GP300) is a prototype array of the GRAND experiment, designed to validate the technique of autonomous radio-detection of astroparticles by detecting cosmic rays with energies between 1017-1018.5 eV. This observation will further enable the study of the Galactic-to-extragalactic source transition region. Between November 2024 and May 2025, 46 out of 300 antennas have been operational and collecting data stably. We present here our cosmic-ray search pipeline, which involves several filtering steps: (1) coincidence search for signals triggering multiple antennas within a time window, (2) directional reconstruction of events, (3) exclusion of clustered (in time and space) noise events, (4) polarization cut, (5) selection based on the size of the footprint, and (6) other less mature cuts in this preliminary stage, including visual cuts. The efficiency of the pipeline is evaluated and applied to the first batch of data, yielding a set of cosmic-ray candidate events, which we present.
@@ 2 Experimental setup and dataset
GP300 consists of 300 antennas spread over an area of 200 km2 situated in the Gobi desert, near Dunhuang, China (lying at N, E). As of July 2025, 65 antennas with an infill step of 577 m have been deployed and are operational, collecting data stably [2]. Each Detection Unit (DU) operates with a sampling frequency of 500 MHz and an analog-to-digital (ADC) converter digitizes signals across three distinct channels, each constructed from a dipole arm of the antenna, along a direction ( channel along North-South, channel along West-East, and channel along the vertical axis). The ADC data, recorded over a duration of 2 s in a frequency range of 50 to 200 MHz, is then converted to voltage units. The GRAND experiment handles data in three formats: Monitoring Data (MD), continuous monitoring of the background radio environments to identify and filter out noise; Unit Data (UD), which records signals from individual DUs that pass a local trigger at DU level based on 6 parameters, including two amplitude thresholds set at 5 and 3 sigma above background noise, ensuring that only transient radio pulses are captured [6, 7]; Coincidence Data (CD), data from multiple (4) antennas triggered within a 10 s time window, providing a more accurate identification and reconstruction of the signal. The collected data is processed via the GRANDlib software [5] in an offline treatment.
@@ 2.1 Data processing: cosmic-ray search pipeline
The pipeline for cosmic-ray search is directly applied to CD events. They are subject to two types of noise: transient noise and stationary background noise. Transient noise induces events and is targeted for batch removal due to its high occurrence; for example, 70 % of all events in the period originate from an electric transformer in the northeast. Stationary background noise, on the other hand, appears within the baseline of the trace but does not independently trigger DUs. For this pipeline, we aim to adopt a conservative approach, prioritizing purity over efficiency in our cosmic-ray identification process.
@@ 2.2 Clustering Cut
The GP300 dataset is affected by interference from commercial aircraft hovering near the array and regular emissions from an electric transformer located northeast of the site. This interference results in bursts of noise events that can be excluded from the analysis. Given that cosmic rays produce short radio pulses ( ns) and thus cannot result in multiple CD events in our data, a clustering algorithm is well-suited for this task.
The algorithm is based on the spatial proximity of events using Planar Wave Front (PWF) arrival direction reconstruction, following the method developed in [8]. Multiple events occurring within the same angular and time window are designated as a cluster and consequently excluded from the dataset. The angular distance between events and is defined as the minimum between and . The temporal distance between and is defined as . Fig. 2 (left) presents the distribution of angular and time distances for consecutive events in our dataset. By varying the angular and time cut parameters below which the events are suppressed, the figure demonstrates the extent of data reduction achieved with different parameter choices. In this study, a conservative approach is adopted, and the parameters of and s are retained, resulting in a cut of 79 % of the CD data, discriminated against as background noise, as seen in Fig. 5 (right).
@@ 2.4 Systematic quality cuts
To ensure high-quality events, several additional quality cuts are implemented in the cosmic-ray search pipeline. They are independently applied to the data, and the order of the processes does not matter. We detail below all the additional cuts performed on the data. The performances of each cut, in terms of percentage of background events removed are presented in the right-hand table in Fig. 5. The cuts based on the antenna number (footprint size) and zenith angle (inclination of the shower) are built on physical grounds: GP300 focuses on specific cosmic-ray parameter ranges, that favor specific ranges in footprint size and zenith angle. The PWF error cut is more artificial as poorly reconstructed events could still be cosmic rays. However, in the current stage, events that cannot be reconstructed consistently with this PWF procedure are not usable (an error based on spherical wavefront is being implemented and tested). The last two cuts (SNR and RMS) are implemented at this stage of data commissioning in order to simplify noise discrimination and help clean events emerge from the sample. They are however to be removed and/or refined for the next stage of the pipeline development, to process the next batch of data with the 65 operational antennas.
@@
Zenith. Given its antenna sparsity, GP300 specializes in detecting highly inclined air showers with zenith angle . The final cut removes events with a reconstructed arrival direction in zenith angle. In addition, reconstruction of near-horizon event is difficult, hence for now we remove as well events with zenith angle . Results are shown in Fig. 5 (right).
@@ 2.5 Manual quality cuts
Trace visual cut. This visual cut is based on the shape of the signal trace in each antenna. Cosmic rays are characterized by short-duration pulses (between 50 and 100 ns) and low-frequency signals.
@@
After passing all cuts, the candidates are reconstructed using three different methods: lateral distribution function [12] from deconvolved electric field reconstruction [3], angular distribution function (ADF) directly from voltage traces and electric field [13, 14], and Graph Neural Networks directly from voltage traces [15]. The majority of the reconstructions confirm the detection of cosmic rays in the expected energy range. These reconstruction methods serve as a basis for an additional quality cut, using multiple criteria: DUs with reconstructed electric fields ( [12] and [13]), a reconstructed energy eV, and a chi-squared value associated with the ADF method lower than 25, as described in [13].
@@ 3 First batch of cosmic-ray candidates
| Cut Parameters | Cut efficiency | ||||
| Clustering |
79 % | |||
| Polarization | 22 % | ||||
| Nb of DUs | 50 % | ||||
| Zenith | 56 % | ||||
| PWF error | 31 % | ||||
| SNR | 58 % | ||||
| RMS |
85 % | |||